OpenStudy (anonymous):

solve the system by using a matrix equation 7x-4y=4 10x-6y=2

OpenStudy (anonymous):

solve the system by using a matrix equation 7x-4y=4 The result is -12

OpenStudy (anonymous):

10x-6y=2 The result is y=5x/3-1/3 y=1/3(5x-1) 2(5x-3y)=2 10x-6y-2=0

OpenStudy (anonymous):

y=-27/41

OpenStudy (anonymous):

and x=-8/41 ...

OpenStudy (anonymous):

lol...just checking if you are awake...correct answer is y=13 and x=8